Greenhab Report – November 14th

Crew 215 GreenHab Report 14-11-2019
GreenHab Officer: Guy Murphy
Environmental control: Door left open for ventilation 12 midday then reclosed 3pm as the day was overcast.
Average temperatures: –
Hours of supplemental light: Automatic system
Daily water usage for crops: 16 litres (2 full blue watering cans).
Daily water usage for research and/or other purposes:
Water in Blue Tank –
Time(s) of watering for crops: 8am, 4:45pm.
Changes to crops: 2 tomato species, mixed pepper, white onion, cucumber seed planted. Dead dry material was broken up for possible future use as mulch.

SEASONAL PLANTING HISTORY
Established plants introduced 03/11/2019

[3x] strawberry plants (Everbearing), in medium metal planter
[1x] spicy orange thyme (thymus fragrantissimus), small terracotta pot
[1x] lemon balm scented geranium (pelargonium x Melissinum), small terracotta pot
[1x] Mint Mentha – Berries & Cream, small terracotta pot
[1x] Pineapple Mint – mentha suaveolens variegated, small terracotta pot
[1x] Ice plant – Delosperma ‘jewel of the desert garnet’, small terracotta pot
[1x] Sempervivum tectorum assorted Hen & Chicks, small terracotta pot
[1x] Rosemary ‘Barbecue’ (Rosmarinus officinalis ‘Barbecue’), large plastic pot
[1x] Rosemary ‘Spice Island’ Rosmarinus officinalis ‘Spice Island’, large plastic pot
[1x] mint, grapefruit, large plastic pot
[2x] small Aloe Vera, in medium metal planter, frost damaged (removed 1311219)

Seeds Planted 04/11/2019 (Large metal planting tub)

Greek Oregano – Origanum heracleoticum
Chives – Allium schoenoprasum
Sweet Marjoram – Origanum majorana
Lemon Balm – Melissa officinalis
Sage – Salvia Officinalis
Thai Basil – Siam Queen
German/Winter Thyme – Thymus vulgaris

Seeds Planted 10/11/2019 (Large metal planting tub)

Coriander – Coriandrum Sativum

Seeds Planted in 12/11/2019 (2 smaller planting tubs)

Carrots ‘Purple Sun F1’
Radish ‘icicle’

Seeds Planted 13/11/2019
Perpetual Spinach, Rocket, Spinach ‘Lakeside F1’in medium metal planter.
Tomato (Chocolate Cherry) and Tomato Pole Roma in small pots for transplantation.

Seeds Planted 14/11/2019
Tomato (Chocolate Cherry) and Tomato Pole Roma in small pots for transplantation.
Onion (Evergreen White Bunching)
Cucumbers (Patio Snackers F1) in hanging pot
Pepper (Mini Bell Color Mix) in small pots for transplantation.

Other seeds stored in the Greenhab are identified as –

Rocket ‘Arugula Rocket Salad ‘Rocky’ (Diplotaxis tenuifolia), 1/2 packet
Spinach ‘Lakeside F1’, 1/2 packet
Carrots ‘Purple Sun F1’ – 1/2 packet
Radish ‘icicle’ – 1/2 packet
Snow Pea (Oregon Giant)
Cucumbers (Patio Snackers F1), 1/2 packet
Onion (Evergreen White Bunching), 1/2 packet
Pepper (Mini Bell Color Mix), 1/2 packet
Perpetual Spinach (Green Leaf Chard), 1/2 packet

Partly used packed of seed have been stored in the provided tin to protect from vermin. This will also allow repropagation later in the season to stagger harvest times and allow resowing if crops are lost.

25 white Narcissus bulbs (Paperwhites ‘Ziva’)

Harvest: None

Support/supplies needed: N/A

EVA Report – November 14th

Crew 215 EVA Report 14-Nov-2019
EVA # 05
Author: Jennifer Lane
Purpose of EVA: Continuing with our sweep for micrometeorites.
Start time: 0915
End time: 1145
Narrative: Returning to Micrometeorite site 5 (MM5), todays crew managed to sweep 30 of
our pre-created grid squares (30m2) for micrometeorites using the magnet on a stick
method. These samples are all bagged and labelled and will later be analyzed to
determine if any of the magnetic particles picked up, are in fact micrometeorites.
Destination: South on Cow Dung Road to the Micrometeorite site, and then back on Cow
Dung Road North to return to the hab.
Coordinates (use UTM NAD27 CONUS): 518473E 4248960N
Participants: Guy Murphy, Steve Whitfield and Larissa Wilson
Road(s) and routes per MDRS Map: South on Cow Dung Road and then on foot 75m to the
site.
Mode of travel: Spirit Rover and Curiosity Rover

Science Report – November 14th

Science Report 14 November 2019
SOL: 4
Crew 215 – Expedition Boomerang Crew
Submitted by Science Officer Andrew Wheeler
Science Dome Operations: Gypsum samples retrieved were cleaned and weighed before
commencing grinding prior to experimentation. These samples were joined by
micrometeorite samples which were processed and packaged for shipment.
Green Hab: Green Hab operations have continued. See Green Hab report.
RAM Operations: N/A
EVA: The fourth EVA for science saw us returning to the MM5 (518462E 4248949N) site
where 30 micrometeorite samples were collected. 40% of the grid has now been sampled.
See EVA Report.
EVA Suit Maintenance: Nominal. See Operations Report.
Psycho-social Stressors Monitoring: Data collection continues through the hexoskin
monitoring vests. The simulated emergency retrieval of the engineer has been postponed,
subject to approval, to 15Nov2019 at the conclusion of his engineering EVA.
Additional activities: Hab operations were disrupted due to ongoing construction.

Operations Report – November 14th

Crew 215 Operations Report 14-Nov-2019
SOL: 4
Name of person filing report: Shane Usher
Non-nominal systems: N/A
Notes on non-nominal systems: N/A
Generator: run
Hours run: 14h 00 min
From what time last night: approx. 1730
To what time this morning: 0730
List any additional daytime hours when the generator was run: N/A
Solar— SOC % (Before generator is run at night): 65%
Diesel Reading – 50% –
Station Propane Reading – 80%.
Ethanol Free Gasoline: N/A
Water (loft tank): 36 gallons
Water Meter: 1457861 units
Water (static tank): 420
Static to Loft Pump used – no
Water in Green Hab: 380 gallons, nominally 20 L (5.7 gallons) used today
Water in Science Dome: 0 gallons
Toilet tank emptied: no
Deimos rover used: not at hab
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: –
Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: –
Spirit rover used: yes
Hours: 110.7
Beginning charge: (Before EVA): 100%
Ending charge: (On return from EVA, before recharging): 73%
Currently charging: yes.
Opportunity rover used: no
Hours: 66.2
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes
Curiosity rover used: yes
Hours: 116.9
Beginning charge: (Before EVA) : 100%
Ending charge: (On return from EVA, before charging): 84%
Currently charging: yes
Notes on rovers:
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): no
Reason for use: N/A
Oil Added? no
ATV Fuel Used: 0 Gals
# Hours the ATVs were used today: 0
Notes on ATVs: Assigned to the director
HabCar used and why, where? With DIRECTOR
CrewCar used and why, where? At Grand Junction awaiting Crew 216
General notes and comments: N/A
Summary of internet: Functional.
Summary of suits and radios:
Summary of Hab operations: Kitchen benchtop being installed.
Summary of GreenHab operations: Atmospheric equalization to the green hab was carried
out during the day, plants were watered (20L). The heater automatically came on at 5pm.
Summary of Science Dome operations: Washed and crushed gypsum samples, Re-bagged
micrometeorite.
Summary of RAM operations: Nominal
Summary of any observatory issues: Nominal
Summary of health and safety issues: All crew healthy.
Questions, concerns and requests to Mission Support: Opportunity had overheated only
and is now operational.

Summary Report – November 14th

Crew 215 Sol Summary Report 14-NOV-2019
Sol: 4
Summary Title: Making Progress
Author’s name: Andrew Wheeler
Mission Status: Active
Sol Activity Summary: After ascertaining that Curiosity rover was operational (it was) our approved EVA for this Sol saw us returning to the micrometeorite sample grid in the stream wash at MM5 (518462E 4248949N). There, Guy, Larissa and Steve collected 30 samples before returning to the hab for lunch. 40% of the plots have now been sampled. During this time, gypsum samples collected from the White Moon location (517724E 4254500N) were begun to be cleaned and weighed in preparation for experimentation. This continued through the afternoon where samples from the micrometeorite grid were added to those being processed and packaged. Despite a rehearsal, construction in the hab caused a postponement of the disabled Marstronaut evacuation during the evening engineering EVA and, consequently, a ‘picnic’ dinner was held in the Green Hab.
Look Ahead Plan: Planning an EVA to MM5 (518462E 4248949N) to resume collection of micrometeorite samples and a demonstration of returning a disabled astronaut into the hab during the afternoon engineering EVA that had been postponed from the previous sol. See EVA requests. Continue monitoring of psycho-social stressors.
Anomalies in work: N/A.
Weather: Marginally below freezing overnight, light winds, clouded over during the day.
Crew Physical Status: Healthy
EVA: Micrometeorite sampling EVA. See EVA report
Reports to be file: Sol Summary Report, Operations Report, Science Report, GreenHab Report, EVA Request, EVA Report, Journalist Report, Daily Photos.
Support Requested: N/A

Operations Report – November 12th

Crew 215 Operations Report 12-Nov-2019
SOL: 1
Name of person filing report: Shane Usher
Non-nominal systems: Opportunity, Suit 5 Helmet
Notes on non-nominal systems: Opportunity brakes do not work, Suit 10 is repaired, Suit 5 Helmet rivetts falling apart, Suit 2 Battery non-charging, tested on multiple suits and will replace battery.
Generator: run
Hours run: 12h
From what time last night: 1730
To what time this morning: 0730
List any additional daytime hours when the generator was run: N/A
Solar— SOC % (Before the generator is run at night): 73%
Diesel Reading – 53%
Station Propane Reading – 73% today.
Ethanol Free Gasoline: N/A
Water (loft tank): 33 gallons
Water Meter: 1457225 units
Water (static tank): 450 gallons
Static to Loft Pump used – no
Water in Green Hab: 293.5 gallons, 24 L (6.5 gallons) used today
Water in Science Dome: 0 gallons
Toilet tank emptied: yes
Deimos rover used: not at the hab
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: –
Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: –
Spirit rover used: yes
Hours: 110.5
Beginning charge: (Before EVA): 62%
Ending charge: (On return from EVA, before recharging): 100%
Currently charging: yes.
Opportunity rover used: no
Hours: 65.6
Beginning charge: 100%
Ending charge: 100%
Currently charging: no
Curiosity rover used: yes
Hours: 116.1
Beginning charge: (Before EVA) : 62%
Ending charge: (On return from EVA, before charging): 100%
Currently charging: yes
Notes on rovers: no
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): no
Reason for use: N/A
Oil Added? no
ATV Fuel Used: 0 Gals
# Hours the ATVs were used today: 0
Notes on ATVs: Assigned to the director
HabCar used and why, where? With DIRECTOR
CrewCar used and why, where? At Grand Junction awaiting Crew 216
General notes and comments: N/A
Summary of internet: Functional
Summary of suits and radios: Suit 2 needs replacement battery to be changed tomorrow, Suit 5 Helmet falling apart, Suit 10 fuse replaced.
Summary of Hab operations: All systems nominal apart from; Stateroom5, Edison screw lightbulb Max 13W required; Fire alarm test button not working in Hab upper level; Fire alarm test button disintegrated in GreenHab; Numerous powerpoints not functional on ground floor of hab. Water pump is very loud.
Summary of GreenHab operations: Atmospheric equalization to the green hab was carried out during the day, plants were watered (24L). The heater automatically came on at 5pm.
Summary of Science Dome operations: All systems nominal.
Summary of RAM operations: Hammers used and returned, 2 fuses taken, electrical tape taken and returned
Summary of any observatory issues: Nominal
Summary of health and safety issues: All crew healthy.
Questions, concerns and requests to Mission Support: Opportunity is now usable.

Research Report – November 12th

Science Report 12 November 2019
SOL: 2
Crew 215 – Expedition Boomerang Crew
Submitted by Science Officer Andrew Wheeler
Science Dome Operations: The first 10 samples from the new micrometeorite sampling grid at MM5 (518462E 4248949N) were received, processed and prepared for dispatch. Previously collected gypsum samples were retrieved from storage in preparation for the ISRU cement/concrete manufacture experiments.
Green Hab: Green Hab operations have continued. See Green Hab report.
RAM Operations: N/A
EVA: The second EVA for science saw us returning to MM5 (518462E 4248949N) to complete setting out the micrometeorite sampling grid and collect the first row of 10 samples. See EVA Report.
EVA Suit Maintenance: Suit 10 has had the blown fuse replaced and is ready to be returned to use. Suit 2 is still being evaluated. See Operations Report.
Additional activities: Acoustic volumes were recorded. Habitat layout was evaluated. Psycho-social stressors continued to be monitored.

EVA Report – November 12th

Crew 215 EVA Report 12 Nov 2019
EVA # 02
Author: Jennifer Lane
Purpose of EVA: Completed deployment of Micrometeorite grid before beginning sweeping
grid squares, as well as exploring local geology.
Start time: 0945
End time: 1220
Narrative: First priority was to finish off placing the strings at site MM5 to divide
the 10×10 m grid into 1 m squares. This was accomplished fairly quickly, so using the magnet on a fishing rod type configuration, sweeping began. After each square the bag is sealed to store the micrometeorites from that 1 m square and labelled before going onto the next square. Row ‘J’ was able to be fully swept and bagged. With some time left, we travelled further south along Cow Dung Road to road 1101 to explore local geology in the search for aggregate suitable for making cement before returning to the hab with some agate to look at under the microscope. Destination: South on Cow Dung Road to the Micrometeorite site, and then back on Cow Dung Road down to road 1101, before returning to the hab.
Coordinates (use UTM NAD27 CONUS): 518473E 4248960N and 519500E 4247300N
Participants: Commander Andrew Wheeler, Shane Usher and Steve Whitfield
Road(s) and routes per MDRS Map: Cow Dung Road and road 1101.
Mode of travel: Spirit & Curiosity Rover

Journalist Report – November 12th

Tues 12 Nov Sol 2
by Guy Murphy
Shane Usher and Steve Whitfield undertook their first EVAs today, on a morning excursion led by our geologist and Mission Commander, Andrew Wheeler. Larissa and Jennifer helped them get ready. The process starts with the donning of flight suits and hiking boots. The astronauts then select and test radios, and put on headsets. They are helped into their single or double piece space suits, which have internal ventilation fans that need to be activated. Once ready, the crew stand in the intermediate airlock space as a 5 minute simulated decompression occurs – a patient wait for those on each side of the airlock door. The explorers then step through the outer airlock door and onto the dusty surface to begin work.
As Greenhab Officer for the mission, I spent spent some of the day working in the outside of the main Hab. The Greenhab is a freestanding greenhouse connected by a simulated pressurised tunnel which is large enough to partially complement the diets of crew with vegetables at peak harvest time. Utah is very cold in winter, but during the day the cloudless skies quickly warm the Greenhab interior. Lights and heaters come on at night.
The operational season for the MDRS has just begun, so we are still at the planting stage. Seeds for kitchen herbs have germinated. Larger plants are in terra cotta pots. Strawberries have begun flowering. Today I planted carrot and radish seed, and soon I will sow tomatoes, capsium and rocket. Given our diet of shelf storable dehydrated food, fresh greens are a luxury on Mars. While labour intensive, it is worth the effort of cultivation. Tending the plants is an enjoyable, relaxing activity. The presence of a Greenhab Officer in each crew should ensure continuity in plant care over the operational season of the campus.
We started lunch with Larissa’s entree of vegetarian pulled pork. (Ingredients included shredded banana skins.) Steve helped prepare a pesto pasta for dinner. Soon crew will be enjoying pasta with pesto made from Greenhab grown Basil.

GreenHab Report – November 12th

Crew 215 GreenHab Report 12-11-2019
GreenHab Officer: Guy Murphy
Environmental control: Door left open for ventilation 12 midday then reclosed 5pm.
Average temperatures: –
Hours of supplemental light: Automatic system
Daily water usage for crops: 24 litres (3 full blue watering cans).
Daily water usage for research and/or other purposes:
Water in Blue Tank – tank was completely filled during afternoon.
Time(s) of watering for crops: 10am, 2pm.
Changes to crops: Carrots, radishes planted.
Established plants introduced to Greenhab 03/11/2019

[3x] strawberry plants (Everbearing), in medium metal planter
[2x] small Aloe Vera, in medium metal planter, frost damaged
[1x] spicy orange thyme (thymus fragantissimus), small terra cotta pot
[1x] lemon balm scented geranium (pelargonium x Melissinum), small terra cotta pot
[1x] Mint Mentha – Berries & Cream, small terra cotta pot
[1x] Pineapple Mint – mentha suaveolens variegated, small terra cotta pot
[1x] Ice plant – Delospermo ‘jewel of the desert garnet’, small terra cotta pot
[1x] Sempervium textorum assorted Hen & Chicks, small terra cotta pot
[1x] Rosemary ‘Barbecue’ (Rosmarinus officinalis ‘Barbecue’), large plastic pot
[1x] Rosemary ‘Spice Island’ Rosmarinus officinalis ‘Spice Island’, large plastic pot
[1x] mint, grapefruit, large plastic pot

Seeds Planted in Greenhab 04/11/2019 (Large metal planting tub)

Greek Oregano – Origanum heracleoticum
Chives – Allium schoenoprasum
Sweet Marjoram – Origanum majorana
Lemon Balm – Melissa officialnis
Sage – Savia Officinalis
Thai Basil – Siam Queen
German/Winter Thyme – Thymus vulgaris

Seeds Planted in Greenhab 10/11/2019 (Large metal planting tub)

Corinader – Coriandrum Sativum

Seeds Planted in Greenhab 12/11/2019 (2 smaller planting tubs)

Carrots ‘Purple Sun F1’
Radish ‘icicle’

Other seeds stored in the Greenhab but not yet planted are identified as –

Onion (Evergreen White Bunching)
Rocket ‘Arugula Rocket Salad ‘Rocky’ (Diplotaxis tenuifolia)
Spinach ‘Lakeside F1’
Tomato Pole Roma ‘San Marzano’ (Solanum lycopersicum)
Carrots ‘Purple Sun F1’ – 1/2 packet left
Radish ‘icicle’ – 1/2 packet left
Snow Pea (Oregon Giant)
Cucumbers (Patio Snackers F1)
Pepper (Mini Bell Color Mix)
Tomato (Chocolate Cherry)
Perpetual Spinach (Green Leaf Chard)

25 white Narcissus bulbs (Paperwhites ‘Ziva’)

Harvest: None

Support/supplies needed: N/A